# From Local to Global: A Graph RAG Approach to Query-Focused Summarization

### v1

@@ 2.5 Graph Communities $\rightarrow$ Community Summaries
The next step is to create report-like summaries of each community in the Leiden hierarchy, using a method designed to scale to very large datasets. These summaries are independently useful in their own right as a way to understand the global structure and semantics of the dataset, and may themselves be used to make sense of a corpus in the absence of a question. For example, a user may scan through community summaries at one level looking for general themes of interest, then follow links to the reports at the lower level that provide more details for each of the subtopics. Here, however, we focus on their utility as part of a graph-based index used for answering global queries.
Community summaries are generated in the following way:
@@
  • Leaf-level communities. The element summaries of a leaf-level community (nodes, edges, covariates) are prioritized and then iteratively added to the LLM context window until the token limit is reached. The prioritization is as follows: for each community edge in decreasing order of combined source and target node degree (i.e., overall prominance), add descriptions of the source node, target node, linked covariates, and the edge itself.
@@
  • Higher-level communities. If all element summaries fit within the token limit of the context window, proceed as for leaf-level communities and summarize all element summaries within the community. Otherwise, rank sub-communities in decreasing order of element summary tokens and iteratively substitute sub-community summaries (shorter) for their associated element summaries (longer) until fit within the context window is achieved.
2.6 Community Summaries $\rightarrow$ Community Answers $\rightarrow$ Global Answer
@@
  • Map community answers. Generate intermediate answers in parallel, one for each chunk. The LLM is also asked to generate a score between 0-100 indicating how helpful the generated answer is in answering the target question. Answers with score 0 are filtered out.
@@
3 Evaluation
3.1 Datasets
We selected two datasets in the one million token range, each equivalent to about 10 novels of text and representative of the kind of corpora that users may encounter in their real world activities:
  • Podcast transcripts. Compiled transcripts of podcast conversations between Kevin Scott, Microsoft CTO, and other technology leaders (Behind the Tech, Scott,2024). Size: 1669 $\times$ 600-token text chunks, with 100-token overlaps between chunks ($\sim$1 million tokens).
  • News articles. Benchmark dataset comprising news articles published from September 2013 to December 2023 in a range of categories, including entertainment, business, sports, technology, health, and science (MultiHop-RAG; Tang and Yang,2024). Size: 3197 $\times$ 600-token text chunks, with 100-token overlaps between chunks ($\sim$1.7 million tokens).
3.2 Queries
Many benchmark datasets for open-domain question answering exist, including HotPotQA (Yang et al.,, 2018), MultiHop-RAG (Tang and Yang,, 2024), and MT-Bench (Zheng et al.,, 2024). However, the associated question sets target explicit fact retrieval rather than summarization for the purpose of data sensemaking, i.e., the process though which people inspect, engage with, and contextualize data within the broader scope of real-world activities (Koesten et al.,, 2021). Similarly, methods for extracting latent summarization queries from source texts also exist (Xu and Lapata,, 2021), but such extracted questions can target details that betray prior knowledge of the texts.
To evaluate the effectiveness of RAG systems for more global sensemaking tasks, we need questions that convey only a high-level understanding of dataset contents, and not the details of specific texts. We used an activity-centered approach to automate the generation of such questions: given a short description of a dataset, we asked the LLM to identify $N$ potential users and $N$ tasks per user, then for each (user, task) combination, we asked the LLM to generate $N$ questions that require understanding of the entire corpus. For our evaluation, a value of $N$ = 5 resulted in 125 test questions per dataset. Table 1 shows example questions for each of the two evaluation datasets.
@@ 5. How can educators use the dataset to highlight the importance of health literacy?
3.3 Conditions
We compare six different conditions in our analysis, including Graph RAG using four levels of graph communities (C0, C1, C2, C3), a text summarization method applying our map-reduce approach directly to source texts (TS), and a naïve “semantic search” RAG approach (SS):
  • C1. Uses high-level community summaries to answer queries. These are sub-communities of C0, if present, otherwise C0 communities projected down.
  • C2. Uses intermediate-level community summaries to answer queries. These are sub-communities of C1, if present, otherwise C1 communities projected down.
  • C3. Uses low-level community summaries (greatest in number) to answer queries. These are sub-communities of C2, if present, otherwise C2 communities projected down.
  • TS. The same method as in subsection 2.6, except source texts (rather than community summaries) are shuffled and chunked for the map-reduce summarization stages.
  • SS. An implementation of naïve RAG in which text chunks are retrieved and added to the available context window until the specified token limit is reached.
The graph index supporting conditions C0-C3 was created using our generic prompts for entity and relationship extraction only, with entity types and few-shot examples tailored to the domain of the data. The graph indexing process used a context window size of 600 tokens with 1 gleaning for the Podcast dataset and 0 gleanings for the News dataset.
Graph RAG Overview of Prominent Public Figures in Entertainment The entertainment industry is vast and diverse, encompassing film, television, music, sports, and digital media. Certain public figures stand out due to their significant contributions and influence across these sectors. The following summary highlights key individuals who are repeatedly mentioned in various entertainment articles, reflecting their impact and presence within the industry. Actors and Directors […] Public Figures in Controversy […] Musicians and Executives […] Athletes and Coaches […] Influencers and Entrepreneurs […] The repeated mention of these figures in entertainment articles signifies their ongoing relevance and the public’s interest in their work. Their influence spans across various aspects of entertainment, from shaping cultural narratives in film and television to driving trends in music and digital media. These individuals not only contribute to their respective fields but also influence the broader cultural landscape, often becoming central figures in social discussions and public discourse.
LLM Decision Comprehensiveness: Winner=1 (Graph RAG)
Answer 1 is better because it provides a more comprehensive and detailed list of public figures from a wider range of entertainment sectors, including film, television, music, sports, gaming, and digital media. It also includes specific examples of their contributions and the impact they have on their respective fields, as well as mentions of controversies and their implications. Answer 2, while detailed in its coverage of a few individuals, is limited to a smaller number of public figures and focuses primarily on their personal lives and relationships rather than a broad spectrum of their professional influence across the entertainment industry.
Diversity: Winner=1 (Graph RAG)
Answer 1 is better because it provides a more varied and rich response by covering a wide range of public figures from different sectors of the entertainment industry, including film, television, music, sports, gaming, and digital media. It offers insights into the contributions and influence of these figures, as well as controversies and their impact on public discourse. The answer also cites specific data sources for each mentioned figure, indicating a diverse range of evidence to support the claims. In contrast, Answer 2 focuses on a smaller group of public figures, primarily from the music industry and sports, and relies heavily on a single source for data, which makes it less diverse in perspectives and insights.
Empowerment: Winner=1 (Graph RAG)
Answer 1 is better because it provides a comprehensive and structured overview of public figures across various sectors of the entertainment industry, including film, television, music, sports, and digital media. It lists multiple individuals, providing specific examples of their contributions and the context in which they are mentioned in entertainment articles, along with references to data reports for each claim. This approach helps the reader understand the breadth of the topic and make informed judgments without being misled. In contrast, Answer 2 focuses on a smaller group of public figures and primarily discusses their personal lives and relationships, which may not provide as broad an understanding of the topic. While Answer 2 also cites sources, it does not match the depth and variety of Answer 1.
Directness: Winner=2 (Naïve RAG)
Answer 2 is better because it directly lists specific public figures who are repeatedly mentioned across various entertainment articles, such as Taylor Swift, Travis Kelce, Britney Spears, and Justin Timberlake, and provides concise explanations for their frequent mentions. Answer 1, while comprehensive, includes a lot of detailed information about various figures in different sectors of entertainment, which, while informative, does not directly answer the question with the same level of conciseness and specificity as Answer 2.
Table 2: Example question for the News article dataset, with generated answers from Graph RAG (C2) and Naïve RAG, as well as LLM-generated assessments.
3.4 Metrics
LLMs have been shown to be good evaluators of natural language generation, achieving state-of-the-art or competitive results compared against human judgements (Wang et al., 2023a, ; Zheng et al.,, 2024). While this approach can generate reference-based metrics when gold standard answers are known, it is also capable of measuring the qualities of generated texts (e.g., fluency) in a reference-free style (Wang et al., 2023a, ) as well as in head-to-head comparison of competing outputs (LLM-as-a-judge, Zheng et al.,2024). LLMs have also shown promise at evaluating the performance of conventional RAG systems, automatically evaluating qualities like context relevance, faithfulness, and answer relevance (RAGAS, Es et al.,2023).
Given the multi-stage nature of our Graph RAG mechanism, the multiple conditions we wanted to compare, and the lack of gold standard answers to our activity-based sensemaking questions, we decided to adopt a head-to-head comparison approach using an LLM evaluator. We selected three target metrics capturing qualities that are desirable for sensemaking activities, as well as a control metric (directness) used as a indicator of validity. Since directness is effectively in opposition to comprehensiveness and diversity, we would not expect any method to win across all four metrics.
@@ 3.4 Metrics
Our head-to-head measures computed using an LLM evaluator are as follows:
@@
  • Empowerment. How well does the answer help the reader understand and make informed judgements about the topic?
  • Directness. How specifically and clearly does the answer address the question?
For our evaluation, the LLM is provided with the question, target metric, and a pair of answers, and asked to assess which answer is better according to the metric, as well as why. It returns the winner if one exists, otherwise a tie if they are fundamentally similar and the differences are negligible. To account for the stochasticity of LLMs, we run each comparison five times and use mean scores. Table 2 shows an example of LLM-generated assessment.
3.5 Configuration
3.6 Results
@@
The indexing process resulted in a graph consisting of 8564 nodes and 20691 edges for the Podcast dataset, and a larger graph of 15754 nodes and 19520 edges for the News dataset. Table 3 shows the number of community summaries at different levels of each graph community hierarchy.
@@
Global approaches vs. naïve RAG. As shown in Figure 4, global approaches consistently outperformed the naïve RAG (SS) approach in both comprehensiveness and diversity metrics across datasets. Specifically, global approaches achieved comprehensiveness win rates between 72-83% for Podcast transcripts and 72-80% for News articles, while diversity win rates ranged from 75-82% and 62-71% respectively. Our use of directness as a validity test also achieved the expected results, i.e., that naïve RAG produces the most direct responses across all comparisons.
Community summaries vs. source texts. When comparing community summaries to source texts using Graph RAG, community summaries generally provided a small but consistent improvement in answer comprehensiveness and diversity, except for root-level summaries. Intermediate-level summaries in the Podcast dataset and low-level community summaries in the News dataset achieved comprehensiveness win rates of 57% and 64%, respectively. Diversity win rates were 57% for Podcast intermediate-level summaries and 60% for News low-level community summaries. Table 3 also illustrates the scalability advantages of Graph RAG compared to source text summarization: for low-level community summaries (C3), Graph RAG required 26-33% fewer context tokens, while for root-level community summaries (C0), it required over 97% fewer tokens. For a modest drop in performance compared with other global methods, root-level Graph RAG offers a highly efficient method for the iterative question answering that characterizes sensemaking activity, while retaining advantages in comprehensiveness (72% win rate) and diversity (62% win rate) over naïve RAG.
@@
Empowerment. Empowerment comparisons showed mixed results for both global approaches versus naïve RAG (SS) and Graph RAG approaches versus source text summarization (TS). Ad-hoc LLM use to analyze LLM reasoning for this measure indicated that the ability to provide specific examples, quotes, and citations was judged to be key to helping users reach an informed understanding. Tuning element extraction prompts may help to retain more of these details in the Graph RAG index.
4 Related Work
4.1 RAG Approaches and Systems
When using LLMs, RAG involves first retrieving relevant information from external data sources, then adding this information to the context window of the LLM along with the original query (Ram et al.,, 2023). Naïve RAG approaches (Gao et al.,, 2023) do this by converting documents to text, splitting text into chunks, and embedding these chunks into a vector space in which similar positions represent similar semantics. Queries are then embedded into the same vector space, with the text chunks of the nearest k vectors used as context. More advanced variations exist, but all solve the problem of what to do when an external dataset of interest exceeds the LLM’s context window.
Advanced RAG systems include pre-retrieval, retrieval, post-retrieval strategies designed to overcome the drawbacks of Naïve RAG, while Modular RAG systems include patterns for iterative and dynamic cycles of interleaved retrieval and generation (Gao et al.,, 2023). Our implementation of Graph RAG incorporates multiple concepts related to other systems. For example, our community summaries are a kind of self-memory (Selfmem, Cheng et al.,2024) for generation-augmented retrieval (GAR, Mao et al.,2020) that facilitates future generation cycles, while our parallel generation of community answers from these summaries is a kind of iterative (Iter-RetGen, Shao et al.,2023) or federated (FeB4RAG, Wang et al.,2024) retrieval-generation strategy. Other systems have also combined these concepts for multi-document summarization (CAiRE-COVID, Su et al.,2020) and multi-hop question answering (ITRG, Feng et al.,2023; IR-CoT, Trivedi et al.,2022; DSP,  Khattab et al.,2022). Our use of a hierarchical index and summarization also bears resemblance to further approaches, such as generating a hierarchical index of text chunks by clustering the vectors of text embeddings (RAPTOR,  Sarthi et al.,2024) or generating a “tree of clarifications” to answer multiple interpretations of ambiguous questions (Kim et al.,, 2023). However, none of these iterative or hierarchical approaches use the kind of self-generated graph index that enables Graph RAG.
4.2 Graphs and LLMs
Use of graphs in connection with LLMs and RAG is a developing research area, with multiple directions already established. These include using LLMs for knowledge graph creation (Trajanoska et al.,, 2023) and completion (Yao et al.,, 2023), as well as for the extraction of causal graphs (Ban et al.,, 2023; Zhang et al.,, 2024) from source texts. They also include forms of advanced RAG (Gao et al.,, 2023) where the index is a knowledge graph (KAPING, Baek et al.,2023), where subsets of the graph structure (G-Retriever,  He et al.,2024) or derived graph metrics (Graph-ToolFormer, Zhang,2023) are the objects of enquiry, where narrative outputs are strongly grounded in the facts of retrieved subgraphs (SURGE, Kang et al.,2023), where retrieved event-plot subgraphs are serialized using narrative templates (FABULA, Ranade and Joshi,2023), and where the system supports both creation and traversal of text-relationship graphs for multi-hop question answering (Wang et al., 2023b, ). In terms of open-source software, a variety a graph databases are supported by both the LangChain (LangChain,, 2024) and LlamaIndex (LlamaIndex,, 2024) libraries, while a more general class of graph-based RAG applications is also emerging, including systems that can create and reason over knowledge graphs in both Neo4J (NaLLM,  Neo4J,2024) and NebulaGraph (GraphRAG, NebulaGraph,2024) formats. Unlike our Graph RAG approach, however, none of these systems use the natural modularity of graphs to partition data for global summarization.

### v2

@@ 3.1.5 Graph Communities $\rightarrow$ Community Summaries
The next step creates report-like summaries of each community in the community hierarchy, using a method designed to scale to very large datasets. These summaries are independently useful as a way to understand the global structure and semantics of the dataset, and may themselves be used to make sense of a corpus in the absence of a specific query. For example, a user may scan through community summaries at one level looking for general themes of interest, then read linked reports at a lower level that provide additional details for each subtopic. Here, however, we focus on their utility as part of a graph-based index used for answering global queries.
GraphRAG generates community summaries by adding various element summaries (for nodes, edges, and related claims) to a community summary template. Community summaries from lower-level communities are used to generate summaries for higher-level communities as follows:
@@
  • Leaf-level communities. The element summaries of a leaf-level community are prioritized and then iteratively added to the LLM context window until the token limit is reached. The prioritization is as follows: for each community edge in decreasing order of combined source and target node degree (i.e., overall prominence), add descriptions of the source node, target node, the edge itself, and related claims.
@@
  • Higher-level communities. If all element summaries fit within the token limit of the context window, proceed as for leaf-level communities and summarize all element summaries within the community. Otherwise, rank sub-communities in decreasing order of element summary tokens and iteratively substitute sub-community summaries (shorter) for their associated element summaries (longer) until they fit within the context window.
3.1.6 Community Summaries $\rightarrow$ Community Answers $\rightarrow$ Global Answer
@@
  • Map community answers. Intermediate answers are generated in parallel. The LLM is also asked to generate a score between 0-100 indicating how helpful the generated answer is in answering the target question. Answers with score 0 are filtered out.
@@
3.2 Global Sensemaking Question Generation
To evaluate the effectiveness of RAG systems for global sensemaking tasks, we use an LLM to generate a set of corpus-specific questions designed to asses high-level understanding of a given corpus, without requiring retrieval of specific low-level facts. Instead, given a high-level description of a corpus and its purposes, the LLM is prompted to generate personas of hypothetical users of the RAG system. For each hypothetical user, the LLM is then prompted to specify tasks that this user would use the RAG system to complete. Finally, for each combination of user and task, the LLM is prompted to generate questions that require understanding of the entire corpus. Algorithm 1 describes the approach.
Algorithm 1: Prompting Procedure for Question Generation
1:Input: Description of a corpus, number of users $K$, number of tasks per user $N$, number of questions per (user, task) combination $M$.
2:Output: A set of $K*N*M$ high-level questions requiring global understanding of the corpus.
3:procedure GenerateQuestions
4:  Based on the corpus description, prompt the LLM to:
  1. 1.
    Describe personas of $K$ potential users of the dataset.
  2. 2.
    For each user, identify $N$ tasks relevant to the user.
  3. 3.
    Specific to each user & task pair, generate $M$ high-level questions that:
    • Require understanding of the entire corpus.
    • Do not require retrieval of specific low-level facts.
5:  Collect the generated questions to produce $K*N*M$ test questions for the dataset.
6:end procedure
For our evaluation, we set $K=M=N=5$ for a total of 125 test questions per dataset. Table 1 shows example questions for each of the two evaluation datasets.
@@ 5. How can educators use the dataset to highlight the importance of health literacy?
3.3 Criteria for Evaluating Global Sensemaking
Given the lack of gold standard answers to our activity-based sensemaking questions, we adopt the head-to-head comparison approach using an LLM evaluator that judges relative performance according to specific criteria. We designed three target criteria capturing qualities that are desirable for global sensemaking activities.
@@ 3.3 Criteria for Evaluating Global Sensemaking
Appendix F shows the prompts for our head-to-head measures computed using an LLM evaluator, summarized as:
@@
  • Empowerment. How well does the answer help the reader understand and make informed judgments about the topic?
Furthermore, we use a “control criterion” called Directness that answers “How specifically and clearly does the answer address the question?”. In plain terms, directness evaluates the concision of an answer in a generic sense that applies to any generated LLM summarization. We include it to behave as a reference against which we can judge the soundness of results for the other criteria. Since directness is effectively in opposition to comprehensiveness and diversity, we would not expect any method to win across all four criteria.
In our evaluations, the LLM is provided with the question, the generated answers from two competing systems, and prompted to compare the two answers according to the criterion before giving a final judgment of which answer is preferred. The LLM either indicates a winner; or, it returns a tie if they are fundamentally similar. To account for the inherent stochasticity of LLM generation, we run each comparison with multiple replicates and average the results across replicates and questions. An illustration of LLM assessment for answers to a sample question can be found in Appendix D.
4 Analysis
4.1 Experiment 1
4.1.1 Datasets
We selected two datasets in the one million token range, each representative of corpora that users may encounter in their real-world activities:
Podcast transcripts. Public transcripts of Behind the Tech with Kevin Scott, a podcast featuring conversations between Microsoft CTO Kevin Scott and various thought leaders in science and technology  (Scott,, 2024). This corpus was divided into 1669 $\times$ 600-token text chunks, with 100-token overlaps between chunks ($\sim$1 million tokens).
News articles. A benchmark dataset comprised of news articles published from September 2013 to December 2023 in a range of categories, including entertainment, business, sports, technology, health, and science  (Tang and Yang,, 2024). The corpus is divided into 3197 $\times$ 600-token text chunks, with 100-token overlaps between chunks ($\sim$1.7 million tokens).
4.1.2 Conditions
We compared six conditions including GraphRAG at four different graph community levels (C0, C1, C2, C3), a text summarization method that applies our map-reduce approach directly to source texts (TS), and a vector RAG “semantic search” approach (SS):
  • C1. Uses high-level community summaries to answer queries. These are sub-communities of C0, if present, otherwise C0 communities projected downwards.
  • C2. Uses intermediate-level community summaries to answer queries. These are sub-communities of C1, if present, otherwise C1 communities projected downwards.
  • C3. Uses low-level community summaries (greatest in number) to answer queries. These are sub-communities of C2, if present, otherwise C2 communities projected downwards.
  • TS. The same method as in Section 3.1.6, except source texts (rather than community summaries) are shuffled and chunked for the map-reduce summarization stages.
  • SS. An implementation of vector RAG in which text chunks are retrieved and added to the available context window until the specified token limit is reached.
The graph index supporting conditions C0-C3 was created using our generic prompts for entity and relationship extraction, with entity types and few-shot examples tailored to the domain of the data.
4.1.3 Configuration
We used a fixed context window size of 8k tokens for generating community summaries, community answers, and global answers (explained in Appendix C). Graph indexing with a 600 token window (explained in Section A.2) took 281 minutes for the Podcast dataset, running on a virtual machine (16GB RAM, Intel(R) Xeon(R) Platinum 8171M CPU @ 2.60GHz) and using a public OpenAI endpoint for gpt-4-turbo (2M TPM, 10k RPM).
We implemented Leiden community detection using the graspologic library (Chung et al.,, 2019). The prompts used to generate the graph index and global answers can be found in Appendix E, while the prompts used to evaluate LLM responses against our criteria can be found in Appendix F. A full statistical analysis of the results presented in the next section can be found in Appendix G.
4.2 Experiment 2
To validate the comprehensiveness and diversity results from Experiment 1, we implemented claim-based measures of these qualities. We use the definition of a factual claim from Ni et al., (2024), which is “a statement that explicitly presents some verifiable facts.” For example, the sentence “California and New York implemented incentives for renewable energy adoption, highlighting the broader importance of sustainability in policy decisions” contains two factual claims: (1) California implemented incentives for renewable energy adoption, and (2) New York implemented incentives for renewable energy adoption.
To extract factual claims, we used Claimify (Metropolitansky and Larson,, 2025), an LLM-based method that identifies sentences in an answer containing at least one factual claim, then decomposes these sentences into simple, self-contained factual claims. We applied Claimify to the answers generated under the conditions from Experiment 1. After removing duplicate claims from each answer, we extracted 47,075 unique claims, with an average of 31 claims per answer.
We defined two metrics, with higher values indicating better performance:
  1. 1.
    Comprehensiveness: Measured as the average number of claims extracted from the answers generated under each condition.
  2. 2.
    Diversity: Measured by clustering the claims for each answer and calculating the average number of clusters.
[…]
5 Results
5.1 Experiment 1
@@
The indexing process resulted in a graph consisting of 8,564 nodes and 20,691 edges for the Podcast dataset, and a larger graph of 15,754 nodes and 19,520 edges for the News dataset. Table 2 shows the number of community summaries at different levels of each graph community hierarchy.
@@
Global approaches vs. vector RAG. As shown in Figure 2 and Table 6, global approaches significantly outperformed conventional vector RAG (SS) in both comprehensiveness and diversity criteria across datasets. Specifically, global approaches achieved comprehensiveness win rates between 72-83% (p<.001) for Podcast transcripts and 72-80% (p<.001) for News articles, while diversity win rates ranged from 75-82% (p<.001) and 62-71% (p<.01) respectively. Our use of directness as a validity test confirmed that vector RAG produces the most direct responses across all comparisons.
@@
Empowerment. Empowerment comparisons showed mixed results for both global approaches versus vector RAG (SS) and GraphRAG approaches versus source text summarization (TS). Using an LLM to analyze LLM reasoning for this measure indicated that the ability to provide specific examples, quotes, and citations was judged to be key to helping users reach an informed understanding. Tuning element extraction prompts may help to retain more of these details in the GraphRAG index.
Community summaries vs. source texts. When comparing community summaries to source texts using GraphRAG, community summaries generally provided a small but consistent improvement in answer comprehensiveness and diversity, except for root-level summaries. Intermediate-level summaries in the Podcast dataset and low-level community summaries in the News dataset achieved comprehensiveness win rates of 57% (p<.001) and 64% (p<.001), respectively. Diversity win rates were 57% (p=.036) for Podcast intermediate-level summaries and 60% (p<.001) for News low-level community summaries. Table 2 also illustrates the scalability advantages of GraphRAG compared to source text summarization: for low-level community summaries (C3), GraphRAG required 26-33% fewer context tokens, while for root-level community summaries (C0), it required over 97% fewer tokens. For a modest drop in performance compared with other global methods, root-level GraphRAG offers a highly efficient method for the iterative question answering that characterizes sensemaking activity, while retaining advantages in comprehensiveness (72% win rate) and diversity (62% win rate) over vector RAG.
5.2 Experiment 2
Table 3 shows the results for the average number of extracted claims (i.e., the claim-based measure of comprehensiveness) per condition. For both the News and Podcast datasets, all global search conditions (C0-C3) and source text summarization (TS) had greater comprehensiveness than vector RAG (SS). The differences were statistically significant (p<.05) in all cases. These findings align with the LLM-based win rates from Experiment 1.
Table 4 contains the results for the average number of clusters, the claim-based measure of diversity. For the Podcast dataset, all global search conditions had significantly greater diversity than SS across all distance thresholds (p<.05), consistent with the win rates observed in Experiment 1. For the News dataset, however, only C0 significantly outperformed SS across all distance thresholds (p<.05). While C1-C3 also achieved higher average cluster counts than SS, the differences were statistically significant only at certain distance thresholds. In Experiment 1, all global search conditions significantly outperformed SS in the News dataset – not just C0. However, the differences in mean diversity scores between SS and the global search conditions were smaller for the News dataset than for the Podcast dataset, aligning directionally with the claim-based results.
For both comprehensiveness and diversity, across both datasets, there were no statistically significant differences observed among the global search conditions or between global search and TS.
Finally, for each pairwise comparison in Experiment 1, we tested whether the answer preferred by the LLM aligned with the winner based on the claim-based metrics. Since each pairwise comparison in Experiment 1 was performed five times, while the claim-based metrics provided only one outcome per comparison, we aggregated the Experiment 1 results into a single label using majority voting. For example, if C0 won over SS in three out of five judgments for comprehensiveness on a given question, C0 was labeled the winner and SS the loser. However, if C0 won twice, SS won once, and they tied twice, then there was no majority outcome, so the final label was a tie.
We found that exact ties were rare for the claim-based metrics. One possible solution is to define a tie based on a threshold (e.g., the absolute difference between the claim-based results for condition A and condition B must be less than or equal to $x$). However, we observed that the results were sensitive to the choice of threshold. As a result, we focused on cases where the aggregated LLM label was not a tie, representing 33% and 39% of pairwise comparisons for comprehensiveness and diversity, respectively. In these cases, the aggregated LLM label matched the claim-based label in 78% of pairwise comparisons for comprehensiveness and 69-70% for diversity (across all distance thresholds), indicating moderately strong alignment.
Table 3: Average number of extracted claims, reported by condition and dataset type. Bolded values represent the highest score in each column.
Condition Average Number of Claims
News Articles Podcast Transcripts
C0 34.18 32.21
C1 32.50 32.20
C2 31.62 32.46
C3 33.14 32.28
TS 32.89 31.39
SS 25.23 26.50
Table 4: Average number of clusters across different distance thresholds, reported by condition and dataset type. Bolded values represent the highest score in each row.
Dataset Distance Threshold Average Number of Clusters
C0 C1 C2 C3 TS SS
News Articles 0.5 23.42 21.85 21.90 22.13 21.80 17.92
0.6 21.65 20.38 20.30 20.52 20.13 16.78
0.7 20.19 19.06 19.03 19.13 18.62 15.80
0.8 18.86 17.78 17.82 17.79 17.30 14.80
Podcast Transcripts 0.5 23.16 22.62 22.52 21.93 21.14 18.55
0.6 21.65 21.33 21.21 20.62 19.70 17.39
0.7 20.41 20.04 19.79 19.22 18.08 16.28
0.8 19.26 18.77 18.46 17.89 16.66 15.07